OpenStudy (anonymous):

The pair of figures is similar. Find x. If necessary, round to the nearest tenth

OpenStudy (anonymous):



OpenStudy (anonymous):

similar, therefore the ratio of the length to width are the same. \[\frac{ width_1 }{ width_2 } = \frac{ length_1 }{length_2 }\] \[\frac{ 3}{ 8 } = \frac{ x }{ 11 }\]

OpenStudy (anonymous):

so its 33?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

4.1

OpenStudy (anonymous):

33/8 = 4.125